# PINN-MEP: Continuous Neural Representations for Minimum-Energy Path Discovery in Molecular Systems

### v5

@@ A.2 Connection between Doob’s Lagrangian and MEP Optimization
We elaborate on the theoretical connections between our energy-based optimization approach and the framework of Doob’s h-transform for sampling transition paths (Du et al., 2024). This comparison is particularly relevant as both methods employ neural networks to represent molecular state transitions—either as a minimum-energy path in our case or the full transition path ensemble in theirs. Notably, the parametrization of the mean trajectory in Du et al. resembles our MEP representation, differing primarily in our use of blending functions and spatial embeddings for the neural network component.
@@
This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape. Intuitively, it measures the amount of "effort"/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
@@
This action functional describes a Wasserstein Lagrangian flow as formalized in (Neklyudov et al., 2023) and (Neklyudov et al., 2024). In this framework, one can learn stochastic dynamics from samples, thereby learning a process that defines a time-dependent density evolving from initial to final states.
@@
The key difference with our PINN-MEP approach is that we minimize the action of a single deterministic path rather than a full distribution of paths. Mathematically, this corresponds to working with a single trajectory through configuration space rather than an ensemble of possible trajectories. This simplification is reflected in our loss function, which contains one fewer expectation than the full Doob’s Lagrangian.
A.2.1 Comparison with Doob’s Lagrangian Method
We provide a direct comparison between our PINN-MEP approach and the method of Du et al. in terms of computational efficiency, as evaluated by the number of force field/energy function calls, and the maximum energies observed in the paths, corresponding to the peak of the potential barrier. To provide a rigorous comparison, we conducted benchmarks using experimental conditions identical to those reported by Du et al. for alanine dipeptide: AMBER14 force field, 300K temperature. Table 1 presents these results alongside those reported in Du et al. (2024) and their recent follow-up work Lee et al. (2025). The only exception is that we do not evaluate trajectory probability, as the assumption of optimization over a non-physical path length makes the calculation of this quantity impossible.
Method | States | # Evaluations | Max Energy | MinMax Energy |
MCMC variable length | CV | 21.02M | 740.70 695.79 | 52.37 |
MCMC* | CV | 1.29B | 288.46 128.31 | 60.52 |
MCMC variable length | relaxed | 187.54M | 412.65 334.70 | 26.97 |
MCMC | relaxed | 10B | N/A | N/A |
MCMC variable length | exact | 10B | N/A | N/A |
MCMC | exact | 10B | N/A | N/A |
Doob’s Cartesian | exact | 38.40M | 726.40 0.07 | 726.18 |
Doob’s Cartesian, 2 Mixtures | exact | 51.20M | 709.38 162.37 | 513.72 |
Doob’s Cartesian, 5 Mixtures | exact | 51.20M | 541.26 278.20 | 247.96 |
Doob’s Internal | exact | 38.40M | -14.62 0.02 | -14.67 |
Doob’s Internal, 2 Mixtures | exact | 51.20M | -15.38 0.14 | -15.54 |
Doob’s Internal, 5 Mixtures | exact | 51.20M | -15.50 0.31 | -15.95 |
Doob’s Seq2Seq | exact | - | 1505.6 0.45 | 245.05 0.02 |
w/ Temperature Annealing | exact | - | 1583.18 0.3 | 592.13 0.26 |
w/ Interatomic Interpolation | exact | - | 1601.57 0.53 | 3.46 0.03 |
MaxLL | exact | - | 1532.45 | 615 |
w/ Temperature Annealing | exact | - | 1599.03 | 233 |
w/ Interatomic Interpolation | exact | - | 1545.32 | 619 |
PINN-MEP | exact | 15700 | -67.22 3.55 | -70.78 |
PINN-MEP requires approximately 3261× fewer energy function evaluations (15.7K vs. 51.20M) compared to the Doob’s Lagrangian approach while achieving a 4.44 times lower maximum energy barrier. The lower number of required evaluations is likely due to the computation of the loss, which requires sampling their path over both , as we do as well, but also the width of the path. Thus, the tradeoff of sacrificing full ensemble characterisation is expected to come with computational efficiency.
Appendix B Comparison with other related Methods
While our primary benchmark comparison focuses on the Doob’s Lagrangian method due to similarity in task and system setup, examining how our approach relates to other methods for transition path discovery/MEP generation is also informative. Here we present results from two related approaches in the literature, though we emphasize that direct numerical comparisons should be treated with caution due to differences in force fields, optimization objectives, and units used.
Method | # Evaluations | Max Energy (eV) |
NEB | 8063 | 11.0014 |
INR | 3046 | 0.4070 |
Method | CVs | # Evaluations | Runtime / Path |
MCMC (Two-Way Shooting) | No | 1B | 100 hours |
Metadynamics | Yes | 1M | 10 hours |
OM Opt. (Diffusion Model) | No | 10K | 50 min |
Table 2 shows results from Ramakrishnan et al. (2025) comparing Nudged Elastic Band (NEB) with their Implicit Neural Representation (INR) approach for MEP discovery. Both our PINN-MEP approach and their INR method utilize lightweight, preexisting molecular dynamics force fields, though we employ different such force fields. Their approach requires a more complex loss function that evaluates both energy and forces along with path parameterization and its gradient (path velocity), but needs fewer force field evaluations than our method. It is important to note that the "Max Energy" reported in Table 2 is measured in electron volts (eV) and appears to be adjusted for ground state energy reference, which makes direct energy comparisons with our method challenging. The difference in force fields used further complicates any direct energy value comparisons between methods.
These tables corroborate our findings in the Doob’s Lagrangian comparison: methods optimizing for a single MEP consistently require orders of magnitude fewer force field evaluations than approaches that characterize the full transition path ensemble.
B.1 Ablation Studies
@@
This section presents a series of ablation studies examining the impact of various components of our PINN-MEP approach. These experiments help validate our design choices and provide insights into the importance of different architectural and training elements. We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method. For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the energy, and the mean number of iterations to reach the lowest maximum energy. We run each setting in the ablations 15 times for a maximum of 25,000 iterations. We use a batch size of 1, where each batch contains 32 values of the progress coordinate for most cases. The exception is when we evaluate the chain-of-states-like regularization, which minimizes the difference between network outputs for adjacent values of ; in this case, the effective batch size is 32, since there are no interactions between different network outputs otherwise.
Due to the nature of physical force fields, slight deviations in the path can lead to large changes in energy, leading to similar settings having very different energies. This can limit the ability to compare the performance of different settings. However, we find that the results are still directionally informative.
B.1.1 General Settings for Ablation Studies
@@
Number of iterations: 25,000 per run
Number of independent runs: 15 per parameter setting
Evaluation frequency: Every 100 training steps
Optimizer: Adam with optax default parameters
Learning rate:
@@
Frames per batch: 32 points sampled along the progress coordinate
Spring constant: (no chain-of-states regularization)
Interpolation method: Cubic Hermite for blending functions
@@
Network depth: 6 hidden layers
Width factor: 4 (network width scales as number of atoms)
Embedding dimension: Automatically set to
@@
Activation function: ReLU throughout the network
Weight initialization: Glorot uniform for hidden layers, zero initialization for final layer
Bias initialization: Default for hidden layers, zero initialization for final layer
B.1.2 Embedding Strategy
We compared sinusoidal time embeddings with different maximum periods against simple linear projections of the progress coordinate. Table 4 shows that linear embedding (None) performed best, achieving the lowest mean energy, though sinusoidal embeddings with period also achieved competitive results. We find these results at odds with the need for sinusoidal embeddings in the larger BPTI system. Perhaps the AIB9 system, which is characterized by one large general conformational rearrangement primarily driven by main-chain dihedral angles, does not need the high frequency features that sinusoidal embeddings provide, but which might be essential for other systems, due to the well-known bias of MLPs toward modeling low frequencies in data.
Embedding | Max Period | MeanMax Energy () | MinMax Energy | Mean Iterations |
Linear (None) | – | 591.35 963.72 | -337.36 | 200873 |
Sinusoidal | 1326.13 2339.35 | -337.36 | 210907 | |
Sinusoidal | 1528.18 2429.05 | -129.90 | 226290 | |
Sinusoidal | 1888.66 3604.47 | -337.36 | 187493 | |
Sinusoidal | 2453.53 3507.47 | -337.36 | 217053 |
B.1.3 Spring Regularization
We tested the effect of chain-of-states-like spring constant regularization, as shown in Table 5. Interestingly, we found that small amounts of spring regularization ( or ) provided the best results, with achieving the lowest mean energy overall. This suggests that a minimal amount of explicit path continuity enforcement can be beneficial when combined with our neural representation, though this is not a strict requirement for the method to work. However, adding a spring constant reintroduces intra-batch dependencies, which means that batch sizes should be of a minimum size for this approach to work effectively. With low batch sizes, this could lead to the corner cutting effects commonly observed in traditional chain-of-states methods, which might make scaling the addition of spring regularization difficult for larger systems. Since we are able to reach the same minimum energy without spring regularization, we recommend against its use, especially when scaling to larger systems where lower batch sizes are preferred.
Spring Constant | MeanMax Energy () | MinMax Energy | Mean Iterations |
-178.22 260.09 | -337.36 | 153533 | |
-155.41 286.63 | -337.36 | 101600 | |
803.61 586.95 | 205.74 | 159033 | |
0.0 | 1460.41 2486.05 | -337.36 | 209133 |
10528.4 7394.89 | 5065.46 | 155400 | |
48639.7 12111.7 | 33143.6 | 107267 |
B.1.4 Network Architecture
We investigated the effect of network depth and width on transition path quality. Table 6 shows that an 8-layer network with a width factor of 4 achieved the best mean energy performance, followed closely by a 6-layer network with width factor 6.
Depth | Width Factor | MeanMax Energy () | MinMax Energy | Mean Iterations |
8 | 4 | 25.02 467.23 | -337.36 | 159333 |
6 | 6 | 72.42 458.72 | -337.36 | 210267 |
8 | 6 | 165.91 1021.79 | -337.36 | 219433 |
8 | 2 | 1434.97 3206.06 | -337.36 | 178167 |
6 | 4 | 1460.34 2486.07 | -337.36 | 209067 |
6 | 2 | 3621.8 3198.25 | 217.83 | 231200 |
4 | 6 | 4242.74 3568.51 | 140.81 | 231967 |
4 | 4 | 5130.75 5758.01 | -328.49 | 228133 |
4 | 2 | 16200.5 8032.18 | 159.21 | 210367 |
B.1.5 Interpolation Type
We examined different interpolation methods for the blending functions that combine the neural network output with the endpoint constraints. Table 7 compares linear, cosine, quintic, and cubic Hermite interpolation. Cosine interpolation performed best in terms of mean energy, though all methods were able to achieve the minimum energy of -337.36 in at least one run, except for quintic. The cubic Hermite interpolation showed higher variance in results.
The specific formulas for each interpolation type are as follows:
Linear interpolation:
Cubic Hermite interpolation:
Quintic interpolation:
Cosine interpolation:
where and are the basis functions that ensure endpoint constraints and , and is the blending function that controls the neural network’s contribution along the path.
Interpolation Type | MeanMax Energy () | MinMax Energy | Mean Iterations |
Cosine | 442.51 769.51 | -337.36 | 198720 |
Linear | 602.61 1057.13 | -337.36 | 219147 |
Quintic | 967.16 1288.57 | -281.14 | 210227 |
Cubic Hermite | 1325.49 2339.75 | -337.36 | 212807 |
B.1.6 AIB9 Network Architecture
Below are the main details of the AIB9 MEP training runs, as shown in the main paper.
Model type: Transition MLP with sinusoidal time embedding
Network depth: 6 hidden layers
Network width: 4 width factor (scaled relative to the number of atoms)
Time embedding: Sinusoidal position encoding with maximum period of 1,000,000
Activation function: ReLU
B.1.7 Training Parameters
Optimizer: Adam
@@ B.1.7 Training Parameters
Batch size: 4
Frames per batch: 32
Spring constant (): 0.0
Interpolation method: Linear
All other parameters are identical to the base settings listed in the ablations section B.1.1.
B.2 BPTI Modeling and Computational Details
B.2.1 System Preparation
B.2.2 Explicit Solvent Shell Construction
For the BPTI system, we developed a sophisticated approach to handle explicit water molecules efficiently:
Initial water box generation: Both start and end configurations were solvated with TIP3P water molecules using a padding distance of 2.0 nm and energy minimized for 15,000 steps.
Water shell extraction: Rather than using the entire water box, we extracted spherical shells of water molecules within a 1.4 nm radius of any protein atom for both the start and end configurations. This significantly reduced the system size while maintaining the crucial first hydration shell.
Radius optimization: We employed a binary search algorithm to determine optimal shell radii that would capture an identical number of water molecules in both start and end configurations, ensuring consistent system sizes throughout the transition.
Water molecule matching: We applied the Hungarian algorithm to optimally pair water molecules between the start and end states based on spatial proximity, minimizing the overall displacement of water molecules during the transition, which ensures that the intial guess for the MEP does not have very large energies stemming from overlapping molecules.
B.2.3 Hybrid Neural Network Architecture
For the BPTI system with explicit solvent, we developed a specialized hybrid transition model to handle the larger system size efficiently:
Protein representation: Neural network with 6 layers, a width factor of 4, and sinusoidal time embedding for the protein atoms (58 residues, 892 atoms)
Water representation: Parameterized spline interpolation between matched water molecules
Control points: 7 control points for the spline interpolation
Water weight: 0.1 (relative weight of water molecule contributions to the loss)
Clamp threshold: (minimum/maximum energy clipping threshold)
B.2.4 Training Parameters
The BPTI MEP was optimized using the following training parameters:
Optimizer: Adam
@@ B.2.4 Training Parameters
Total iterations: 30,000
All other parameters are identical to the base settings listed in the ablations section B.1.1.
B.2.5 Handling High-Energy Spikes During Training
Physical force fields present unique optimization challenges due to occasional extreme energy spikes when atoms approach one another too closely. To address this, we implemented adaptive gradient clipping in our training procedure:
where is a clipping threshold (set to in our experiments). We specifically use adaptive gradient clipping rather than a fixed threshold because our hybrid architecture contains two fundamentally different parameter types: neural network weights for protein coordinates and spline control points for water molecules. The adaptive nature of the clipping automatically adjusts to the appropriate scale for each parameter type.
Additionally, we apply different scaling factors to these parameter groups:
with for water-related parameters. This differential treatment further balances the influence of each component during optimization. While this approach does not completely eliminate training instabilities, we find it sufficient for most protein systems when combined with careful initialization and learning rate scheduling.
B.2.6 Force Field Parameters

### v6

@@ E.1 Connection between Doob’s Lagrangian and MEP Optimization
We elaborate on the method of Doob’s h-transform for sampling transition paths (Du et al., 2024). This explanation is relevant as both methods employ neural networks to represent molecular state transitions, either as a minimum-energy path in our case or the complete transition path ensemble in theirs. Notably, the parametrization of the mean trajectory in Du et al. resembles our MEP representation, differing primarily in our use of blending functions and spatial embeddings for the neural network component.
@@
This functional quantifies the cost of controlling a stochastic process to achieve desired endpoint conditions, subject to the system’s energy landscape. Intuitively, it measures the amount of ”effort”/action needed to steer the dynamics from starting state A to target state B, with smaller values indicating more probable transition paths.
@@
This action functional describes a Wasserstein Lagrangian flow as formalized in (Neklyudov et al., 2023) and (Neklyudov et al., 2024). In this framework, one can learn stochastic dynamics from samples, thereby learning a process that defines a time-dependent density evolving from an initial to a final state.
@@
The difference with our approach is that we minimize the action of a single deterministic path rather than a complete distribution of paths. This simplification is reflected in our loss function, which contains one fewer expectation than the full Doob’s Lagrangian.
Appendix F Ablation Studies
@@
This section presents a series of ablation studies examining the impact of various components of our AdaPath approach. These experiments help validate our design choices and provide insights into the importance of different architectural and training elements. We conduct these ablation studies on the AIB9 system because Alanine dipeptide, as a benchmark system, is too simple to provide meaningful insights into the performance of our method. For each ablation study, we report the lowest maximum energy of the path among all training runs, the mean and standard deviation of the arc-length adjusted energy, and the mean number of iterations to reach the lowest maximum energy. We run each setting in the ablations 5 times. We use a batch size of 1, where each batch contains 16 values of the progress coordinate for most cases.
F.1 General Settings for Ablation Studies
@@
Number of iterations: 100,000 per run
Number of independent runs: 5 per parameter setting
Evaluation frequency: Every 2000 training steps
Optimizer: Muon as implemented in Optax.
Learning rate: 1e-2
@@
Frames per batch: 16 points sampled along the progress coordinate
Interpolation method: Linear
@@
Network depth: 6 MLP blocks.
MLP Expansion Factor: 4, meaning that each MLP block, consisting of two linear layers, has four times the hidden dimensionality as input and output dimensionality.
Atom Embedding Dimension: 32.
@@
Activation function: GeLu throughout the network
Conditioning MLP settings: 3 Layers, with a hidden dimension of 1024 and GeLu activations.
F.2 Architecture Comparison: AdaPath vs Standard MLP
To validate our core claim that parameter sharing across atoms provides superior scalability, we compare AdaPath against a standard MLP baseline that directly maps progress coordinate to the complete molecular configuration without parameter sharing. This comparison is conducted across all three molecular systems (Alanine dipeptide, AIB9, and BPTI) to demonstrate scaling behavior.
The standard MLP baseline uses a conventional feedforward architecture with six hidden layers of width determined by 4 times the number of atoms in the system. This was found to be suitable for the two small systems. Intuitively, this aligns well with deep learning’s collective wisdom of having a higher hidden dimension than the input and output dimensions of the network. We only lower this to one time the atom count for the BPTI system, as this already leads to an MLP with 1= billion parameters, as more was not runnable for our hardware. The network takes sinusoidal embeddings of the progress coordinate as input and directly outputs the -dimensional molecular configuration, where is the number of atoms in the system. The standard MLP (TransitionMLP) uses the same interpolation strategy and endpoint constraints as AdaPath but processes all atomic coordinates through shared dense layers rather than using per-atom embeddings with parameter sharing.
All other training parameters remain identical to the base configuration described in Section F.1, with both architectures using the same learning rate of 1e-2 with the Muon optimizer for the two smaller systems and the Nadamw optimizer for BPTI with the learning rate scheduling as outlined in G.
| Architecture | Parameters | Max Energy | MinMax Energy | Mean Energy |
| Standard MLP | 5e4 | |||
| AdaPath | 2e6 |
| Architecture | Parameters | Max Energy | MinMax Energy | Mean Energy |
| Standard MLP | 2e6 | |||
| AdaPath | 2e6 |
| Architecture | Parameters | Max Energy | MinMax Energy | Mean Energy |
| Standard MLP | 1.4e9 | |||
| AdaPath | 3.5e6 |
The results demonstrate AdaPath’s superior performance and scalability across system sizes. The AdaPath’s parameter count is dominated by the progress coordinate MLP that outputs the adaLn parameters. This produces a baseline parameter count that puts the AdaPath architecture above the MLP for small systems, but for larger ones, this gap inverts sharply.
On the alanine dipeptide system, AdaPath achieves significantly lower and more consistent maximum energies (-86.6 vs -68.1 kJ/mol for minimum maximum energy) with substantially reduced variance. The performance gap becomes dramatic on the larger AIB9 system, where the standard MLP achieves a best maximum energy of 2.99e4 kJ/mol compared to AdaPath’s -644 kJ/mol, representing a 30,000+ kJ/mol improvement in path quality.
F.3 Optimizer Comparison: Path Continuity Analysis
To validate our claim that the Muon optimizer maintains better path continuity compared to Adam for small systems, we conduct a comparative analysis using path velocity metrics over training iterations. We train identical AdaPath networks on the AIB9 system using both optimizers with learning rates of 1e-2 for Muon and 1e-4 for Adam, respectively. All other training parameters remain identical to the base configuration.
We track the maximum path velocity over training iterations, computed as , where high velocities indicate rapid changes in molecular configuration that can lead to discontinuous ”teleportation” between states rather than smooth transitions.
| Optimizer | Max Energy | Mean Energy |
| Adam | ||
| Muon |
The Adam-trained system exhibits extremely high path velocities, indicating that it rapidly switches from one state to another without passing through the transition region. This is corroborated by the low mean but high max energies.
F.4 Energy Transformation Analysis
We evaluate the impact of our logarithmic energy transformation on training stability by comparing performance with and without the log transformation applied to the potential energy in the loss function. The logarithmic transformation is applied as where ensures positivity, while the baseline uses the direct energy . All other training parameters remain identical to the base configuration.
| Energy Transform | Max Energy | MinMax Energy | Mean Energy |
| Direct Energy | |||
| Log Transform |
The logarithmic transformation achieves significantly lower energies with reduced variance, demonstrating improved training stability and path quality over direct energy optimization.
F.4.1 Reference Trajectory Generation
To generate reference conformational data, we performed an extended molecular dynamics simulation using OpenMM with the following parameters:
Force field: AMBER ff15ipq-m
Temperature: 500 K (elevated to enhance sampling of conformational transitions)
Integrator: Langevin Middle Integrator with friction coefficient of 1 ps-1
Time step: 1 femtosecond
Non-bonded treatment: No cutoff (NoCutoff method)
Constraints: Hydrogen bonds constrained using the HBonds method
Total simulation time: 900 nanoseconds
Output frequency: Coordinates saved every 1000 steps (1 ps intervals)
The simulation protocol consisted of initial energy minimization followed by velocity assignment at the target temperature, a brief 10 ps equilibration period, and then the production run. The elevated temperature of 500 K was chosen to accelerate conformational sampling and ensure adequate transitions between metastable states during the simulation timeframe.
F.4.2 State Definition and Analysis
To characterize the conformational landscape, we focused on the backbone dihedral angles of the central residues. Specifically, we computed the and dihedral angles for residues at indices [56, 58, 60, 69] and [58, 60, 69, 71], respectively, corresponding to the central amino acids of the peptide chain.
Two metastable states were defined using ellipsoidal boundaries in the - dihedral space:
These elliptical regions capture the distinct conformational basins observed in the free energy landscape projected onto the central residue dihedral angles.
F.4.3 Neural Network Training Setup
For the neural network-based MEP optimization, we employed a different force field setup optimized for differentiable molecular dynamics:
Force field: AMBER ff15ipq-m for protein mimetics, implemented through DMFF
Non-bonded treatment: No cutoff method for consistency with the reference simulation
Simulation box: 11.0 nm cubic box
Neighbor list cutoff: 5.0 nm
F.4.4 Start- and End-state Setup
Start and end configurations for MEP optimization were randomly sampled from the reference trajectory frames that satisfied the respective state definitions. The Kabsch algorithm was employed to optimally align the end configuration to the start configuration, removing translational and rotational degrees of freedom. This alignment procedure ensures that the neural network focuses on learning the internal conformational changes rather than rigid-body motions, leading to more efficient and physically meaningful transition path discovery.
Appendix G BPTI System Preparation
G.1 Explicit Solvent System Preparation
For the BPTI system, we developed an approach to handle explicit water molecules and ions using periodic boundary conditions:
System solvation: Both start and end protein configurations were placed in 3.5 nm cubic periodic boxes and solvated with TIP3P water molecules using OpenMM’s Modeller class. The system included Na+ and Cl- ions at 0.04 M ionic strength to approximate physiological conditions.
Energy minimization: To preserve the reference protein conformations from the D.E. Shaw simulation while allowing solvent relaxation, all protein heavy atoms were fixed by setting their masses to zero during energy minimization. This completely freezes the protein structure while allowing water molecules and ions to relax.
Molecule matching: Since the endpoint configurations were solvated independently, we equalized atom counts by randomly removing excess molecules, then used the Hungarian algorithm to optimally pair water molecules and ions between configurations based on center-of-mass positions. This matching minimizes solvent displacement during transitions.
The resulting system contained approximately 3,500 atoms total in a periodic cubic box, maintaining proper solvation while using standard periodic boundary conditions compatible with conventional MD force fields.
G.2 Training Parameters
The BPTI MEP optimization employed a two-stage training strategy due to spiking loss late during training. This is likely due to sharp minima in large molecular configurations, including water, where small changes in atomic coordinates can produce drastic energy spikes. This necessitated transitioning to more conservative learning parameters in Stage 2 to maintain training stability.
Initial Training, iterations 0-6,000:
@@ G.2 Training Parameters
Gradient clipping threshold:
Fine-tuning, iterations 6,000-12,000:
@@
Gradient clipping threshold:
Architecture and General Training Parameters:
MLP Expansion Factor: 2
Layers: 5
Hidden dimension: 64
Progress embedding dimension: 1024
Conditioning depth: 3 layers
Conditioning hidden size: 1024
Optimizer: NAdam with weight decay (nadamw)
Weight decay:
Beta parameters: , ,
@@
Total iterations: 12,000
Evaluation frequency: Every 1000 iterations
G.3 Force Field Parameters